OpenStudy (samsan9):

A paratrooper jumps from an airplane traveling 120 km/hr at a height of 400 meters? If his parachute must open 200 meters above the ground, how long does he have for the chute to open?

OpenStudy (samsan9):

@fateal

OpenStudy (fateal):

lel my sis had a similar problem like this yesterday....so umm the answer must be half :/ i mean he have to open it half way.... so umm do u have any formla for it? that will help allot

OpenStudy (samsan9):

well the formula for time i use is t=d/v

OpenStudy (samsan9):

distance over velocity

OpenStudy (fateal):

so its going to be T= 400/200? or T= 400/120? sorry i havent took math for 2 years...just find out the answer for both then check ur choices?

OpenStudy (samsan9):

alright thank you so much

OpenStudy (samsan9):

i think i have to divide both but ill check tomorrow

OpenStudy (fateal):

anytime ^^ sorry couldnt help that much...just do it n ya check tomorrow...goo luck :)